OpenStudy (anonymous):

x/4 = x - 2 /3 x = 4 x = 8 x = 2

OpenStudy (anonymous):

the / mark means over x over 4 x - 2 over 3

OpenStudy (kropot72):

x/4 = (x - 2)/3 Multiply both sides by 4 and get: x = (4x - 8)/3 Multiply both sides by 3 and get: 3x = 4x - 8 -x = -8 x = 8